What is the molality of 60.5% by mass of nitric acid (HNO3) solution?

1
Expert's answer
2022-03-13T15:20:56-0400

Solution:

solute = nitric acid (HNO3)

solvent = water (H2O)


Suppose we are given 100 g of an aqueous HNO3 solution

Therefore,

Mass of HNO3 = Mass of solution × w(HNO3) / 100% = (100 g × 60.5%) / (100%) = 60.5 g

Mass of H2O = Mass of solution − Mass of HNO3 = 100 g − 60.5 g = 39.5 g


The molar mass of HNO3 is 63.01 g/mol

Therefore,

Moles of HNO3 = (60.5 g HNO3) × (1 mol HNO3 / 63.01 g HNO3) = 0.960 mol HNO3


Kilograms of H2O = (39.5 g H2O) × (1 kg / 1000 g) = 0.0395 kg H2O


Molality = Moles of solute / Kilograms of solvent

Therefore,

Molality of HNO3 solution = Moles of HNO3 / Kilograms of H2O

Molality of HNO3 solution = (0.960 mol) / (0.0395 kg) = 24.3 mol/kg = 24.3 m

Molality of HNO3 solution = 24.3 m


Answer: The molality of nitric acid (HNO3) solution is 24.3 m